OpenStudy (anonymous):

68 in scientific notation?

OpenStudy (anonymous):

\[6.8 \times 10 ^{-1}\]

OpenStudy (anonymous):

dont just give the answer please its against the coc

OpenStudy (anonymous):

@DanielJ they come on here to get help which means showing them how to do the question not just give answers

OpenStudy (anonymous):

I apologise. In future i will demonstrate, as i have in my other replies, how to reach the answer instead of just supplying it. Scientific notation always has one integer, than a decimal point, followed by \[\times 10^{x}\] which denotes how much the comma has moved. In this case, \[6.8 \times 10 ^{-1}\]the -1 indicates that the comma has been moved one place to the left